# ElasticTok: Adaptive Tokenization for Image and Video

### v1

@@ Appendix B Training Details
The tables below showing trainin details for each of our models. Our Long Video model is trained on a mix of images and video (Batch Split), and each run (e.g. Long Video (2)) is initialized from the previous run (e.g. Long Video (1)). For the discrete (FSQ) model, each block has 4k tokens (256 blocks = 1M tokens), and the continuous (VAE) model has 2k tokens in each block (256 blocks = 512K tokens).
ImageNet Long Video (1) Long Video (2) Long Video (3)
Batch Size $256$ $256$ $256$ $256$
# Blocks $1$ $1$ $2$ $4$
Batch Split (Image / Video) $100\%/0\%$ $50\%/50\%$ $50\%/50\%$ $25\%/75\%$
Total Iterations 200k 200k 80k 50k
Learning Rate $2\times 10^{-4}$ $2\times 10^{-4}$ $2\times 10^{-4}$ $2\times 10^{-4}$
Optimizer AdamW AdamW AdamW AdamW
Weight Decay $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$
Warmup Iterations 10k 10k 5k 2k
Long Video (4) Long Video (5) Long Video (6) Long Video (7)
Batch Size $64$ $16$ $8$ $4$
# Blocks $16$ $64$ $128$ $256$
Batch Split (Image / Video) $25\%/75\%$ $25\%/75\%$ $25\%/75\%$ $25\%/75\%$
Total Iterations 10k 1k 500 200
Learning Rate $2\times 10^{-4}$ $2\times 10^{-4}$ $2\times 10^{-4}$ $2\times 10^{-4}$
Optimizer AdamW AdamW AdamW AdamW
Weight Decay $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$
Warmup Iterations 1k 200 100 25
@@ Appendix D Reconstruction Examples by MSE
We include the reconstructed images at different Mean Squared Error (MSE) loss thresholds in Figure 10 as a reference for how the images appear at various MSE thresholds.
Figure 10: Comparison of reconstructed image quality at varying MSE thresholds. The ground truth (GT) image is displayed on the left, followed by reconstructions at different MSE thresholds (0.04, 0.02, 0.01, 0.007, and 0.004), showing progressive improvement in fidelity as the threshold decreases.

### v2

@@ Appendix B Training Details
The tables below showing trainin details for each of our models. Our Long Video model is trained on a mix of images and video (Batch Split), and each run (e.g. Long Video (2)) is initialized from the previous run (e.g. Long Video (1)). For the discrete (FSQ) model, each block has 4k tokens (256 blocks = 1M tokens), and the continuous (VAE) model has 2k tokens in each block (256 blocks = 512K tokens). ImageNet Long Video (1) Long Video (2) Long Video (3) Batch Size $256$ $256$ $256$ $256$ Blocks $1$ $1$ $2$ $4$ Batch Split (Image / Video) $100\%/0\%$ $50\%/50\%$ $50\%/50\%$ $25\%/75\%$ Total Iterations 200k 200k 80k 50k Learning Rate $2\times 10^{-4}$ $2\times 10^{-4}$ $2\times 10^{-4}$ $2\times 10^{-4}$ Optimizer AdamW AdamW AdamW AdamW Weight Decay $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ Warmup Iterations 10k 10k 5k 2k Long Video (4) Long Video (5) Long Video (6) Long Video (7) Batch Size $64$ $16$ $8$ $4$ Blocks $16$ $64$ $128$ $256$ Batch Split (Image / Video) $25\%/75\%$ $25\%/75\%$ $25\%/75\%$ $25\%/75\%$ Total Iterations 10k 1k 500 200 Learning Rate $2\times 10^{-4}$ $2\times 10^{-4}$ $2\times 10^{-4}$ $2\times 10^{-4}$ Optimizer AdamW AdamW AdamW AdamW Weight Decay $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ Warmup Iterations 1k 200 100 25
@@ Appendix D Reconstruction Examples by MSE
We include the reconstructed images at different Mean Squared Error (MSE) loss thresholds in Figure 10 as a reference for how the images appear at various MSE thresholds. Refer to caption Figure 10: Comparison of reconstructed image quality at varying MSE thresholds. The ground truth (GT) image is displayed on the left, followed by reconstructions at different MSE thresholds (0.04, 0.02, 0.01, 0.007, and 0.004), showing progressive improvement in fidelity as the threshold decreases.
Appendix E Details on Neural Regression Model
To collect training data, we first run inference with the tokenizer for a pre-defined noise target of $0.003$ on 100K sequences to get pairs of image/video and token lengths. To train the neural regression model, we augment the pretrained encoder model with alternating downsampling and transformer blocks. We downsample by a factor of 4 over sequence each time each time using a simple size 4 stride 4 1D conv, and each set of transformer blocks consists of 4 layers. The final output is pooled, and fed through a 2 hidden layer MLP to output a single scalar per block. The regressor is optimized using MSE loss, and regresses to a normalized token length in the range $[-1,1]$.
Appendix F Further Experiments
Table 3: Performance of ElasticTok-VAE on videos. Values in the table show the percentage of reconstructed video blocks that satisfy a given reconstruction threshold. The baseline is a 50% fixed token baseline, and our method uses variable token lengths with an average of 50% token usage over the dataset.
{NiceTabular}
@lccc@ Threshold
$\mathbf{0.001}$ $\mathbf{0.003}$ $\mathbf{0.015}$
Baseline $40\%$ $78\%$ $99\%$
Ours $\mathbf{50\%}$ $\mathbf{88\%}$ $\mathbf{99\%}$
Table F above shows the same metrics as Fig 4 but on our continuous variant ElasticTok-VAE model. Similar to ElasticTok-FSQ, the VAE variant can reliably encode more videos to satisfy a given reconstruction threshold while using the same number of tokens (on average) as a FLOPs identical continuous variant baseline.
Table 4: Quality baseline comparison with ElasticTok-FSQ on videos. Values shown in the table represent the average percentage of tokens used by each model to achieve the same worse-case MSE (compute MSE over each video block, and take the max MSE over the all blocks in the video). Our method uses fewer tokens to acheive the same worse-case reconstruction as the baseline.
{NiceTabular} @llll@ Threshold
0.001 $\mathbf{0.003}$ $\mathbf{0.015}$
Baseline $78\%$ $57\%$ $15\%$
Ours $\mathbf{75\%}$ $\mathbf{37\%}$ $\mathbf{9.4\%}$
Table F above shows running a quality baseline comparison between our method and a baseline. For our method, we encode each evaluation video using a given reconstruction threshold, and compute the average percentage of tokens per frame. For the baseline, we find the needed fixed tokens per frame such that the worse-case reconstruction MSE over the whole video matches the worst-case MSE of ElasticTok’s reconstruction. Intuitively, fixed token baselines perform worse due to allocated a fixed number of tokens per video block, and must optimize for the worst-case scenario. In contrast, ElasticTok can save tokens when the videos transition to simpler to encode scenes while allocating more tokens to more complex scenes.
Table 5: Ablation comparing with and without encoder conditioning on the masked tokens. Values in the table show the percentage of reconstructed video blocks that satisfy a given reconstruction threshold under the constraint that both models use the same average percentage of tokens.
{NiceTabular}
@lccc@ Threshold
0.001 0.003 0.015
With Enc Cond 36% 79.6% 98.5%
Without Enc Cond 33.1% 75.3% 98.4%
Table F shows a comparison of ElasticTok with and without encoder conditioning on the mask. ElasticTok with conditioning slightly outperforms the model without conditioning. As such, it may be more efficient to train the model without encoder conditioning depending on expected size of the downstream models - as this present a trade-off in terms of slightly better encoding quality at the cost of 2x slower inference speed.
Appendix G More Progressive Reconstruction Examples
Refer to caption Figure 11: Progressive reconstruction of a given images as you increasing the number of tokens (left-to-right in the encoding). Images are in raster scan order, and the number of tokens used is in the bottom right of each image (from 256 - 4096 interpolated with 20 values on a log scale)
Refer to caption Figure 12: Progressive reconstruction of a given images as you increasing the number of tokens (left-to-right in the encoding). Images are in raster scan order, and the number of tokens used is in the bottom right of each image (from 256 - 4096 interpolated with 20 values on a log scale) Refer to caption Figure 13: Progressive reconstruction of a given images as you increasing the number of tokens (left-to-right in the encoding). Images are in raster scan order, and the number of tokens used is in the bottom right of each image (from 256 - 4096 interpolated with 20 values on a log scale) Refer to caption Figure 14: Progressive reconstruction of a given images as you increasing the number of tokens (left-to-right in the encoding). Images are in raster scan order, and the number of tokens used is in the bottom right of each image (from 256 - 4096 interpolated with 20 values on a log scale)